OpenStudy (anonymous):

How do you solve ln(x+3) - ln(x+4) = -3

OpenStudy (jamesj):

Remember that: \[ \ln a - \ln b = \ln(a/b) \] Can you do it now?

OpenStudy (anonymous):

yes! logs just confuse me